Question 3: Suppose the real rate is 1.88 percent and the inflation rate is 0.89 percent. What rate would you expect to earn on a Treasury bill? Use exact formula and answer in percent to two decimals.
Answer #1

(1+real rate)=(1+nominal rate)/(1+inflation rate)

(1+0.0188)=(1+nominal rate)/(1+0.0089)

1+nominal rate=(1.0188*1.0089)

nominal rate=(1.0188*1.0089)-1

which is equal to

=2.79%(Approx).